Given the following information for Groto Corp. find the WACC. Assume the company's tax rate is 40%. Bonds: 10,000 9% coupon bonds outstanding, $1,000 par value, 25 years to maturity, selling for 98% of par; the bonds make semi-annual payments. Common shares: 300,000 shares outstanding, selling for $40 per share: the beta is 0.95. Preferred shares: 55,000 shares of 6% preferred stock outstanding, currently selling for $100 per share. 10% market risk premium and 3% risk-free rate.
Answer #1

answer

cost of debt

effective interesr rate = using financial calculator PV=-980 N=25 x 2= 50 PMT=90/2 = 45 FV=1000 Calculate = I/Y

effective interest rate =4.60

cost of debt = 4.6(1-t)

=4.6 (1-0.3)

=4.6 (0.7)

=3.22

cost of equity using CAPM formula = Risk free rate + (beta x market risk premium)

=3+(0.95 x 10)

= 3 + 9.5

=12.5%

cost of preference shares = 6%

market capital = value of bonds + value of equity + value of prefernce

=10000 x 1000 + 300,000 x 40 + 55000 x 100

= 27500000

WACC = cost of debt x percentage debt in cap + cost of equity x percentage equity in cap + cost of prefernce x percentage of prefrence in cap

=(10,000,000/27,500,000) x 3.22% + (12,000,000/27,500,000) x 12.5% + (5,500,000/27,500,000) x 6%

=0.36 x 3.22% +0.44 x 12.5% + 0.2 x 6%

= 1.16 + 5.5 + 1.2

WACC=8.3%